OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

divide by -3 \[x+7=4 or x+7=-4\] solve for x

OpenStudy (anonymous):

divide everything by three?

OpenStudy (anonymous):

no only -12 because you have to isolate x

OpenStudy (anonymous):

you will get 2 answers x+7=4 x+7=-4 solve for x in both of those

OpenStudy (anonymous):

x=3?